OpenStudy (anonymous):

What is the exact value of cos 17pi/8? a) square root of 2+the square root of 2/4 b.) 0.38 c.) 0.99 d.)square root of 2-the square root of 2/4 ***My Answer: D***

OpenStudy (loser66):

Nope

OpenStudy (anonymous):

Dangit hahaha. My second choice would have to be A then

OpenStudy (loser66):

Nope again, ahhahha

OpenStudy (anonymous):

Dangit(lol)...Not the best at trig identities. Sorry

OpenStudy (loser66):

ok, hint 17pi/8 = 16pi/8 + pi/8

OpenStudy (loser66):

and you have the perfect angle 16pi/8 = 2pi

OpenStudy (loser66):

break it out by cos (a+b) =....

OpenStudy (anonymous):

How would I find the a and b?

OpenStudy (anonymous):

Would it just be a= 16pi/8 and b= pi/8?

OpenStudy (loser66):

hey, \(cos (\dfrac{17\pi}{8}=cos (\dfrac{16\pi +\pi}{8}\)

OpenStudy (loser66):

yup

OpenStudy (anonymous):

Oh I see. Ok

OpenStudy (anonymous):

So from there we get it into the form...\[\cos(\frac{ 16\pi }{ 8 })\cos(\frac{ \pi }{ 8 }) + \sin(\frac{ 16\pi }{ 8 })\sin(\frac{ \pi }{ 8 })\]

OpenStudy (anonymous):

I think

OpenStudy (anonymous):

@Loser66 Now this is the part that I have trouble understanding. I do not know where to go from here

OpenStudy (loser66):

whyyyyyyyyy? 16pi/8 = 2pi, right?

OpenStudy (anonymous):

Yes

OpenStudy (loser66):

cos 2pi =1, right?

OpenStudy (loser66):

sin 2pi =0, ok?

OpenStudy (anonymous):

Ok

OpenStudy (loser66):

so, at the end, you just have cos (pi/8 ) =0.99999

OpenStudy (anonymous):

Ohhhhhh ok! That was my problem all along! I would divide pi/8 without the cos! Haha Thank you so much!

OpenStudy (loser66):

ok

OpenStudy (anonymous):

What the... my computer just told me that it was A :(

OpenStudy (loser66):

Facepalm. hehehe... let's get other's help @campbell_st

OpenStudy (mathstudent55):

\(\dfrac{17\pi}{8} = \dfrac{16\pi}{8} + \dfrac{\pi}{8} = 2\pi + \dfrac{\pi}{8}\) \(\cos \dfrac{17\pi}{8} = \cos 2\pi + \dfrac{\pi}{8} = \cos \dfrac{\pi}{8} \) \(\Large \cos \dfrac{\pi}{8} = \cos \left ( {\dfrac{\frac{\pi}{4}}{2} } \right)\) \(\cos \dfrac{\theta}{2} = \sqrt{\dfrac{1}{2}(1 + \cos \theta) }\) \(\Large \cos \dfrac{\pi}{8} = \cos \left ( {\dfrac{\frac{\pi}{4}}{2} } \right) = \sqrt{\dfrac{1}{2}(1 + \cos \dfrac{\pi}{4})}\) \(= \sqrt{\dfrac{1}{2} (1 + \dfrac{\sqrt{2}}{2})} = {\sqrt{\dfrac{1}{2} + \dfrac{\sqrt{2}}{4}}} \) \(= \sqrt{\dfrac{2 + \sqrt 2}{4}} = \dfrac{\sqrt{2 + \sqrt 2 } }{ 2 }\)

OpenStudy (loser66):

WWWWWWWWWWWWoooooah!!! It is much.... more logic than my way. Thank you so much @mathstudent55

OpenStudy (mathstudent55):

You're welcome.

OpenStudy (mathstudent55):

@Loser66 You were trying to use the identity for the cosine of a sum of angles. That method does not work in this problem. That method works if you can break up the angle into two angles whose cosines you know. The problem here is that the cos of pi/8 is not one of the know values. On the other hand, the cos of pi/4 is well known, so using the identity of the cosine of a half angle gives you the result.